OpenStudy (anonymous):

Find the zero's of this quadratic. x^2=4x+8

OpenStudy (anonymous):

\[x ^{2}=4x+8\]

OpenStudy (anonymous):

@jim_thompson5910 what do you get for this one?

jimthompson5910 (jim_thompson5910):

were you able to get any answers at all?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

what did you get

OpenStudy (anonymous):

\[2 \pm -2\sqrt{3}\]

OpenStudy (anonymous):

what did you get?

jimthompson5910 (jim_thompson5910):

nailed it again

jimthompson5910 (jim_thompson5910):

x^2=4x+8 x^2-4x=8 x^2-4x-8 = 0 Use the quadratic formula to solve for x \[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\Large x = \frac{-(-4)\pm\sqrt{(-4)^2-4(1)(-8)}}{2(1)}\] \[\Large x = \frac{4\pm\sqrt{16-(-32)}}{2}\] \[\Large x = \frac{4\pm\sqrt{16+32}}{2}\] \[\Large x = \frac{4\pm\sqrt{48}}{2}\] \[\Large x = \frac{4+\sqrt{48}}{2} \ \text{or} \ x = \frac{4-\sqrt{48}}{2}\] \[\Large x = \frac{4+4\sqrt{3}}{2} \ \text{or} \ x = \frac{4-4\sqrt{3}}{2}\] \[\Large x = 2+2\sqrt{3} \ \text{or} \ x = 2-2\sqrt{3}\] So I'm getting the same thing.

OpenStudy (anonymous):

thanks so much! I did it a little differently but I guess it didn't matter :)

jimthompson5910 (jim_thompson5910):

yeah as long as you get the same final answer, that's all that matters